# MedArabiQ: Benchmarking Large Language Models on Arabic Medical Tasks

### v1

@@ 3.1.6 Q&A with Grammatical Error Correction (GEC)
Since the patient Q&A dataset uses dialectal Arabic, we constructed an additional version with enhanced linguistic quality and consistency. We specifically applied a Grammatical Error Correction (GEC) pipeline tailored for Arabic healthcare texts. Given the morphological complexity and syntactic richness of Arabic, this preprocessing step was essential. First, we used cameltools, an open-source Arabic NLP library, to morphologically disambiguate words and dediacritize text, ensuring uniformity (Obeid et al., 2020). This step was crucial for handling Arabic’s inflectional patterns and preparing the dataset for grammatical correction. Then, we employed an existing fine-tuned BERT-based Arabic Grammatical Error Detection (GED) model (Alhafni et al., 2023) to detect agreement errors, incorrect word order, and missing inflections. Each token was tagged with a grammatical label to facilitate structured corrections. Next, we used the GEC model from the same pipleine (built on mBART), to automatically correct detected errors while preserving semantic integrity. The model was fine-tuned on the QALB-2015, QALB-2014, and ZAEBUC corpora to enhance its performance in grammatical corrections (Mohit et al., 2014; Rozovskaya et al., 2015; Habash and Palfreyman, 2022).
@@ 3.1.7 Q&A with LLM Modifications
Additionally, to mitigate potential model memorization, since some models could have been trained using scraped data of Altibbi, we modified the dataset using an LLM for a more rigorous assessment of reasoning and adaptability (Dong et al., 2024a). In particular, we used GPT-4o to paraphrase our original questions using the prompt, “You are a helpful assistant that paraphrases text while keeping its meaning intact.” This approach ensured that the core medical concepts remained unchanged while reducing the likelihood of models relying on memorized content.
In summary, we constructed seven new datasets using two primary sources (AraMed and past medical exams) via extensive manual verification to build the MedArabiQ benchmark.
@@
tab:benchmark_results Benchmark Datasets Metrics LLaMA Claude GPT-4o Qwen Gemini Multiple Choice Questions Accuracy 26.2 53.5 53.5 38.0 57.5 Fill-in-the-Blank with choices Accuracy 42.4 72.7 70.7 42.40 72.7 Fill-in-the-Blank without choices BERTScore 78.2 83.6 88.2 83.5 88.8 Patient-Doctor Q&A BERTScore 81.2 81.1 84.5 85.2 82.5 Q&A with Grammatical Error Correction BERTScore 85.5 83.7 84.5 85.2 82.5 Q&A with LLM modifications BERTScore 84.9 83.6 84.2 84.9 82.3
Table 1: Summary of performance results across six benchmark datasets. We present the results in terms of accuracy and BERTScore, depending on task, and show best results in bold per row (Note we evaluate LLaMA 3.1, Claude-3.5-Sonnet, GPT-4o, Qwen 2.5, and Gemini 1.5 Pro).
3.2 Models
Benchmarks used to evaluate LLMs often face the challenge of data contamination. Many modern LLMs are trained on large web crawl datasets sourced from the internet, which can inadvertently include benchmark questions. This overlap inflates performance metrics, making such benchmarks unreliable for fair evaluation (Chunyuan et al., 2024; Dong et al., 2024b). Addressing this issue is critical to ensure that benchmark results accurately reflect the true capabilities of LLMs, free from biases introduced by prior exposure to evaluation data.
In addition to categorizing models based on the transparency of their training data, this study further classifies models according to their linguistic coverage. Specifically, we distinguish between models that are broadly multilingual and those that are primarily Arabic-centric multilingual. Although many large language models include training data from numerous languages, exposure to a language does not inherently confer proficiency (Bender et al., 2021). GPT-4, for example, is known to have been trained on a variety of languages (Achiam et al., 2023). By assessing multilingual and Arabic-centric multilingual models separately, we determine how language coverage and specialization influence performance, particularly in the domain of Arabic medical tasks. Table C1 shows the models selected.
3.3 Evaluation
The evaluation of LLMs in Arabic healthcare applications requires a comprehensive framework that balances technical performance with real-world applicability (Sallam and Mousa, 2024). Our framework assesses medical reasoning, decision-making, and dialogue-based question answering (Kanithi et al., 2024; Guo et al., 2023). Multiple choice questions were evaluated based on accuracy, while open-ended questions, such as those in the fill-in-the-blank and patient-doctor Q&A, were assessed using BERTScore to measure semantic alignment (Zhang* et al., 2020).
3.4 Bias Assessment and Mitigation
To ensure that LLMs can be effectively deployed in healthcare and provide meaningful contributions to the field, it is critical to address their potential to replicate human biases. Completely eliminating bias from LLMs is nearly impossible, as the datasets used to train these models are inherently shaped by human judgment, which is itself susceptible to bias. Recognizing this challenge, we developed an evaluation framework to systematically assess model resistance to biases injected into the multiple choice questions, measure susceptibility, and test mitigation strategies. The framework draws on methodologies outlined in recent work (Schmidgall et al., 2024), while adapting key elements to the context of Arabic medical datasets and healthcare applications. These adaptations include culturally relevant bias categories, prompts designed for clinical scenarios, and additional evaluation metrics to align with real-world needs.
  1. 1.
  2. 2.
    Bias Testing: Models are tested with biased variations of prompts. We compute the change in accuracy to assess impact of bias.
  3. 3.
@@ 3.4 Bias Assessment and Mitigation
Figure 3: Performance Samples from Closed and Open-ended Benchmarks. We report model outputs for samples from (a) Multiple Choice Questions and Fill-in-the-Blank, and (b) Patient-Doctor Q&A tasks, illustrating differences in accuracy, response validity, and language consistency across closed and open-source models.
4 Experimental Setup
Evaluation. We perform zero-shot prompting for all models and tasks, adjusting the temperature according to the nature of each benchmark after an initial testing phase. For closed models, we set the temperature to 0.2 for all tasks. For open-source models, we set the temperature to 0 for Q&A and fill-in-the-blank closed tasks to ensure deterministic answers, and 0.4 for open-ended tasks. For BERTScore evaluation, the model used is XLM-RoBERTa-Large, as it was trained on multiple languages, including Arabic, making it more suitable than monolingual models.
Instruction Prompts. Prompt engineering is crucial for evaluating LLMs. In our prompt engineering experiments, we tested both English and Arabic prompts and found that English prompts were generally more effective. However, for AraMed benchmark tasks, Arabic prompts performed better in open-source models. Based on these findings, we used English prompts for all tasks except patient-doctor Q&A, Q&A with GEC, and Q&A with LLM modifications. Supplementary Table D1 lists the prompts used.
Answer Processing. For Q&A, Q&A with bias, and fill-in-the-blank with choices, models generate both the index of the predicted correct option and the full textual answer. However, some models, particularly open-source ones, exhibit spelling inconsistencies. To ensure accuracy, we evaluated only the first character generated after the phrase “The correct letter is:”, comparing it against valid options. If the character does not match a valid choice, then the response is deemed invalid. For open-ended benchmarks, where answers are not constrained to multiple choice formats, we evaluate the full model-generated response.
@@ 4 Experimental Setup
5 Results
We report all results of the experiments performed on all six benchmarks in Table 1. Our results show that no single model outperforms all others across all benchmarks. For closed tasks (Multiple Choice Questions and fill-in-the-blank with and without choices), closed models perform best, as expected. Gemini 1.5 Pro achieves the highest accuracy in three out of six benchmark tasks, although it performs comparably with Claude 3.5 Sonnet on one. For Multiple Choice Questions and fill-in-the-blank with choices, Gemini performs best, with accuracy scores of 57.5 and 72.7, respectively. Claude 3.5 Sonnet matches Gemini’s 72.7 accuracy on the fill-in-the-blank dataset. In open-ended tasks (Patient-Doctor Q&A, Q&A with GEC, and Q&A with LLM Modifications), Qwen performs best, achieving 85.2 on Patient-Doctor Q&A, while LLaMa scores best on AraMed Q&A with GEC with a score of 85.5. However, models do not consistently excel across both task types, as Qwen and LLaMa perform poorly on closed tasks.
@@ 5 Results
Figure 2 (a) compares the performance of open-source and API-based models averaged across all benchmark tasks. Among all models, Gemini 1.5 Pro achieves the highest average benchmark score, followed closely by GPT-4 and Claude 3.5, reinforcing the dominance of proprietary models in NLP tasks. The open-source models, such as Llama 3.1 and Qwen 2.5, demonstrate competitive but more variable performance, with Qwen 2.5 emerging as the strongest among them. The error bars indicate that while closed models exhibit greater stability across tasks, open-source models show higher performance variability, likely due to their reliance on general-purpose pretraining rather than domain-specific fine-tuning. These findings highlight the current dominance of proprietary models while underscoring the potential of open-source models with further fine-tuning and domain adaptation.
@@
Figure 2 (b) provides a task-specific performance breakdown, revealing significant performance variation across multiple choice, fill-in-the-blank, and generative Q&A tasks. Closed models consistently achieve high scores across all benchmarks, particularly excelling in Q&A tasks, including those with GEC and LLM-enhanced responses. This suggests a strong capability for handling complex medical inquiries, a critical requirement in real-world applications. However, the fill-in-the-blank and multiple choice tasks exhibit more performance divergence, with open-source models like Qwen 2.5 and Llama 3.1 lagging behind the leading closed models. The observed discrepancies emphasize the need for fine-tuning on structured medical datasets, particularly for open-source models, to improve accuracy in knowledge-intensive tasks. Figure 3 shows examples of model responses for the closed and open-ended tasks.
@@
Figure 4 explores the impact of bias and its mitigation strategies on model performance. In Figure 4 (a), we compare the accuracy of GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet-20240620 on original questions versus questions embedded with bias across various bias categories. The results reveal that, generally, all models experience a decline in accuracy when bias is introduced, with the extent of decrease varying by bias type. Gemini 1.5 Pro demonstrates a relatively higher resilience to bias, especially in Status Quo Bias and False-Consensus Bias, while Claude 3.5 Sonnet shows more significant performance drops in Self-Diagnosis Bias and Cultural Bias. Additionally, Figure 4 (b) illustrates the effectiveness of different bias mitigation strategies. Notably, all models exhibit improved accuracy with strategies like Few-Shot prompting compared to biased questions without mitigation. Gemini 1.5 Pro consistently outperforms the other models across these strategies, reinforcing its robustness in handling biased inputs.
@@
Figure 4 (c) presents a radar plot summarizing model performance across the employed bias mitigation strategies. The plot shows Gemini’s consistent performance across all strategies, while GPT-4o and Claude 3.5 Sonnet display larger variability across strategies such as Bias Education and One-Shot prompting. These results further highlight the importance of bias mitigation strategies in enhancing model reliability. Additional information on the performance of models by bias category and medical specialty across different bias mitigation strategies in Figures F1 and G1 in the Appendix.
@@ 6 Discussion
Our findings are consistent with prior research on medical LLM benchmarking, where closed-source models outperform open-source models in structured tasks but perform comparably in generative ones. Chen et al. (2025) demonstrated that proprietary models like GPT-4 and Med-PaLM 2 excel in multiple choice and factual retrieval tasks, largely due to stronger pretraining on structured datasets and domain-specific knowledge integration. Similarly, Alonso et al. (2024) found that proprietary models generally achieve higher accuracy in medical question-answering tasks across multiple languages, reinforcing the idea that closed models have better knowledge grounding and factual consistency. Our results support these trends, with Gemini 1.5 Pro and Claude 3.5 Sonnet leading in multiple choice and fill-in-the-blank tasks, suggesting that API-based models are better suited for clinical decision support and structured question-answering tasks.
@@
Figure 4 further illustrates the variability in model performance when exposed to biased content and the effectiveness of different bias mitigation strategies for the top 3 scoring models (GPT-4o, Gemini 1.5 Pro, and Claude 3.5 Sonnet-20240620). All models experience an accuracy drop when bias is introduced, with Claude 3.5 Sonnet showing the most pronounced decline across multiple bias categories. Notably, bias mitigation techniques such as few-shot prompting have shown significant improvement in model performance, especially for Gemini 1.5 Pro, which shows the greatest resilience against bias. Yet, no single strategy consistently mitigates bias across all models and categories, underscoring the complexity of bias in medical NLP and the need for further research.
@@
For generative Q&A tasks, however, traditional automatic evaluation metrics such as BERTScore do not fully capture actual model performance. Although GPT-4 and Claude 3.5 produce responses that are highly relevant and contextually accurate as shown in 3, their longer responses result in lower BERTScores compared to ground truth references. This issue has been previously noted by Liu et al. (2025) who found that metrics like ROUGE and BERTScore struggle to effectively assess medical LLMs due to the inherent complexity of medical diagnoses, where multiple treatment options may exist for a single patient. Our findings reinforce this limitation, as models with lower BERTScores sometimes generate high-quality, informative answers that are penalized for verbosity rather than inaccuracy. This discrepancy highlights the need for more nuanced evaluation approaches, including expert human assessment or task-oriented dialogue evaluations, to better reflect real-world usability in medical consultations.
7 Ethical Considerations
@@
Given the sensitive nature of healthcare applications, ethical considerations are paramount when developing and deploying large language models (LLMs). Our benchmark, MedArabiQ, was carefully constructed from publicly available and anonymized educational resources, ensuring compliance with data privacy standards by excluding personally identifiable patient information. Furthermore, we recognize and discuss several ethical implications associated with clinical deployment of LLMs, including potential risks such as misinformation, unintended amplification of biases (cultural, confirmation, or status quo biases), and limited model interpretability. Addressing these concerns, we emphasize the importance of comprehensive validation through hybrid frameworks combining automated evaluation metrics with clinician expert reviews. We advocate for continuous monitoring, clinician oversight, and clearly defined operational guidelines to mitigate potential harm and biases. By explicitly detailing these ethical considerations, we aim to promote responsible and transparent adoption of LLMs, ultimately contributing to safer and more equitable healthcare solutions.”
8 Limitations and Future Work
@@
While our study provides a comprehensive evaluation of Arabic medical LLMs, there are areas that warrant further exploration. First, there is a possibility for contamination in the datasets. Although the past medical exams are not available in structured digital format and required an extensive effort for digitization and cleaning, we cannot completely rule out the possibility of contamination. That being said, model performance can be improved highlighting the importance of the benchmark datasets. As for AraMed, we explicitly incorporated modifications to test for potential memorization considering that it is a publicly available dataset. To verify the validity of the dataset, preliminary results were obtained in collaboration with medical students to assess the Q&A dataset for factuality, relevance, clarity, and complexity on a scale from 1 to 5. The average scores (Relevance: 4.99, Factuality: 4.97, Accuracy: 4.88, Clarity: 4.89) provide initial evidence of the dataset’s reliability and utility. Extensive evaluation is required in future work as the scope and size of the benchmark is expanded.
@@
Furthermore, our benchmarks are in Modern Standard Arabic (MSA), following standardization practices to ensure consistency. However, this does not account for dialectal variations, which are commonly used in real-world patient-doctor interactions. Future work could explore incorporating dialectal data to enhance model adaptability across diverse Arabic-speaking healthcare contexts. Additionally, while this study focuses on text-based benchmarks as a unimodal foundation, expanding Arabic medical NLP benchmarks to support multimodal AI, such as medical images and lab results, would be a valuable direction for future research.
@@
9 Conclusion
In this work, we introduced the first structured benchmark for evaluating LLMs in Arabic healthcare, addressing a critical gap in Arabic medical NLP. Our benchmark consists of 700 diverse clinical samples, covering both structured medical knowledge assessments and real-world patient-doctor interactions. Beyond Arabic healthcare, our benchmark lays the foundation for developing benchmarks in other medically underserved languages, contributing to the global refinement of medical AI applications.

### v2

@@ 4.1.6 Q&A with Grammatical Error Correction (GEC)
Since the patient-doctor Q&A dataset uses dialectal Arabic, we constructed an additional version with enhanced linguistic quality and consistency. We specifically applied a Grammatical Error Correction (GEC) pipeline tailored for Arabic healthcare texts. Given the morphological complexity and syntactic richness of Arabic, this preprocessing step was essential. First, we used cameltools, an open-source Arabic NLP library, to morphologically disambiguate words and dediacritize text, ensuring uniformity (Obeid et al., 2020). This step was crucial for handling Arabic’s inflectional patterns and preparing the dataset for grammatical correction. Then, we employed an existing fine-tuned BERT-based Arabic Grammatical Error Detection (GED) model (Alhafni et al., 2023) to detect agreement errors, incorrect word order, and missing inflections. Each token was tagged with a grammatical label to facilitate structured corrections. Next, we used the GEC model from the same pipleine, built on mBART, to automatically correct detected errors while preserving semantic integrity. The model was fine-tuned on the QALB-2015, QALB-2014, and ZAEBUC corpora to enhance its performance in grammatical corrections (Mohit et al., 2014; Rozovskaya et al., 2015; Habash and Palfreyman, 2022).
@@ 4.1.7 Q&A with LLM Modifications
Additionally, to mitigate potential model memorization, since some models could have been trained using scraped data of Altibbi, we modified the dataset using an LLM for a more rigorous assessment of reasoning and adaptability (Dong et al., 2024). In particular, we used GPT-4o to paraphrase our original questions using the prompt, “You are a helpful assistant that paraphrases text while keeping its meaning intact.” This approach ensured that the core medical concepts remained unchanged while reducing the likelihood of models relying on memorized content (Zhou et al., 2024).
In summary, we constructed seven new datasets using two primary sources, AraMed and past medical exams, via extensive manual verification to build the MedArabiQ benchmark, all of which are summarized in Table F1 in Appendix F.
4.2 Data Extraction
In the collection of the MCQ dataset in Section 4.1.1 and the fill-in-the-blank questions dataset in Section 4.1.3, we started off by collecting paper-based past exams and lecture notes sourced from a large repository of academic materials hosted on student-led social platforms of regional medical schools. No personally identifiable information or real patient data was included, and thus anonymization was not necessary as our data collection complied with privacy and ethical guidelines. These exams and answers were not readily available in structured digital formats, necessitating a rigorous manual process to ensure clarity and correctness. Given that Arabic medical education is not widely digitized, these exams are not publicly accessible in structured formats. Even if some individual questions exist online, the extensive effort required to compile, format, and structure them into a benchmark dataset significantly reduces the likelihood of contamination. Questions were selected to reflect increasing complexity across different academic years, ensuring that model performance could be assessed at varying levels of medical expertise.
As for the patient-doctor Q&A dataset, our selection process prioritized questions with well-formed queries and meaningful answers, avoiding instances where responses were overly generic (e.g., “Consult a doctor” or “See a specialist”). However, we retained some examples of such cases to reflect real-world user behavior, as patients often seek medical advice even for questions that are better suited for in-person consultations.
Additionally, we maintained proportionality in representation to mirror the dataset’s original structure, particularly for categories like reproductive and sexual health. These categories comprised a significant portion of the dataset and are often underrepresented in Arabic-language medical research despite their importance. Given the sensitivity and cultural taboos associated with these topics, we ensured their inclusion to provide a more realistic and balanced evaluation of the model’s ability to handle diverse medical inquiries. This approach ensures the dataset reflects the varied nature of medical concerns while maintaining its relevance to real-world healthcare scenarios. We selected 100 questions, allocating an equal proportion of samples to each specialty: cardiology, obstetrics and gynecology, surgery, pediatrics, neurology, oncology, endocrinology, dentistry, otholarhyntology, public health, dermatology, primary care, pulmonology, and psychology.
We also incorporated information about the patient, specifically age and gender, when known, into the query. When this information was unknown or erroneous (i.e., the age was an illogical number), it was excluded. Some questions came from an acquaintance of the patient, so no information was known about the patient. The information was usually prepended to the question or inserted after the greeting, in the format of “I am a [man/woman] and I am [x] years old”. The inclusion of this information was useful in incorporating personalization into the case scenarios, making our benchmark dataset better suited to evaluate LLM performance in real-world medical scenarios.
5 Results
We report all results of the experiments performed on six of our benchmarks in Table 1. Our results show that no single model outperforms all others across all benchmarks. For closed tasks (MCQ and fill-in-the-blank with and without choices), closed models perform best, as expected. Gemini 1.5 Pro achieves the highest accuracy in two out of six benchmark tasks, while Deepseek leads on the fill-in-the-blank with choices task with an accuracy of 79.7. For closed tasks, Gemini and Deepseek perform best with accuracy scores of 57.5 and 79.7, respectively. In open-ended tasks, namely patient-doctor Q&A, Q&A with GEC, and Q&A with LLM modifications, Jais performs best, achieving scores of 85.7, 85.6, and 85.5 respectively. However, models do not consistently excel across both task types, as Jais performs poorly on closed tasks. Among open-access models, Jais is the best-performing, followed by LLaMA, which achieves the highest score on Q&A with LLM Modifications.
@@ 5 Results
tab:benchmark_results Benchmark Datasets Metrics Falcon Jais LLaMA Claude GPT-4 Qwen Gemini Deepseek Multiple-Choice Questions Accuracy 20.2 16.0 26.2 53.5 53.5 38.0 57.5 50.5 Fill-in-the-Blank with choices Accuracy 20.2 49.4 42.4 72.7 70.7 42.40 72.7 79.7 Fill-in-the-Blank without choices BERTScore 85.1 86.2 78.2 83.6 88.2 83.5 88.8 87.4 Patient-Doctor Q&A BERTScore 84.9 85.7 81.2 81.1 84.5 85.2 82.5 82.2 Q&A with LLM modifications BERTScore 84.8 85.6 85.5 83.7 84.5 85.2 82.5 82.3 Q&A with Grammatical Error Correction BERTScore 84.4 85.5 84.9 83.6 84.2 84.9 82.3 82.0
Table 1: Summary of performance results across all benchmark datasets. We present the results in terms of accuracy and BERTScore, depending on task, and show best results in bold per row. Due to space constraints, abbreviated model names are used in the table; full names are as follows: Falcon = Falcon3-10B-Instruct, Jais = jais-13b-chat, LLaMA = LLaMA-3.1-8B-Instruct, Claude = claude-3-5-sonnet-20240620, GPT-4 = gpt-4-0613, Qwen = Qwen2.5-7B-Instruct, Gemini = Gemini 1.5 Pro, Deepseek = Deepseek v3.
@@
Figure 3: Performance Samples from Closed and Open-ended Benchmarks. We report model outputs for samples from (a) MCQ and Fill-in-the-Blank, and (b) patient-doctor Q&A tasks, illustrating differences in accuracy, response validity, and language consistency across proprietary high-resource and open-access models. An English version of this figure is provided in Figure G1 of the appendix.
@@
Figure 2 (a) compares the performance of open-access and proprietary high-resource models averaged across all benchmark tasks. Among all models, Gemini 1.5 Pro achieves the highest average benchmark score, followed closely by GPT-4, Deepseek v3 and Claude 3.5, reinforcing the dominance of proprietary high-resource models in NLP tasks. The open-access models, including Llama 3.1, Qwen 2.5, and Falcon, demonstrate competitive but more variable performance, with Qwen 2.5 emerging as the strongest among them. The error bars indicate that while closed models exhibit greater stability across tasks, open-access models show higher performance variability, likely due to their reliance on general-purpose pretraining rather than domain-specific fine-tuning. These findings highlight the current dominance of proprietary high-resource models while underscoring the potential of open-access models with further fine-tuning and domain adaptation.
@@
Figure 2 (b) provides a task-specific performance breakdown, revealing significant performance variation across MCQ, fill-in-the-blank, and generative Q&A tasks. Closed models consistently achieve high scores across all benchmarks, particularly excelling in Q&A tasks, including those with GEC and LLM-modified questions. This suggests a strong capability for handling complex medical inquiries, a critical requirement in real-world applications. However, the fill-in-the-blank and multiple-choice tasks exhibit more performance divergence, with open-access models like Qwen 2.5 and Llama 3.1 lagging behind the leading closed based models. Notably, Jais performs better in certain benchmarks but struggles with structured tasks, suggesting limitations in factual consistency and retrieval-based reasoning. The observed discrepancies emphasize the need for fine-tuning on structured medical datasets, particularly for open-access models, to improve accuracy in knowledge-intensive tasks. Figure 3 shows examples of model responses for the closed and open-ended tasks.
[…]
tab:bias_evaluation_results Model Q&A Q&A with Bias Bias Education One-Shot Mitigation Few-Shot Mitigation Claude 3.5 Sonnet 53.1 52.0 51.0 55.1 52.0 GPT-4 66.3 35.7 42.9 46.9 46.9 Gemini 1.5 Pro 55.1 55.1 57.1 54.1 59.2
Table 2: Comparison of performance by accuracy score without bias, with bias, and with bias mitigation across the three chosen models.
@@
Table 2 and Figure 4 highlight the impact of bias and different mitigation strategies on model performance. In Figure 4 (a), we compare the accuracy of GPT-4, Gemini 1.5 Pro, and Claude 3.5 Sonnet-20240620 on original questions versus questions injected with bias across various bias categories. The results reveal that, generally, all models experience a decline in accuracy when bias is introduced, with the extent of decrease varying by bias type. Across all bias catgories, Gemini 1.5 Pro demonstrates high resilience to bias, while GPT-4 shows more significant performance drops in False-Consensus Bias and Recency Bias. Additionally, Figure 4 (b) illustrates the effectiveness of different bias mitigation strategies. Notably, all models exhibit improved accuracy with strategies like Few-Shot prompting compared to biased questions without mitigation. Gemini 1.5 Pro outperforms the other models with both Bias Education and Few-Shot Mitigation, reinforcing its robustness in handling biased inputs.
@@
Figure 4 (c) presents a radar plot summarizing model performance across the employed bias mitigation strategies. The plot shows Gemini’s consistent performance across all strategies, while GPT-4 and Claude 3.5 Sonnet display larger variability across strategies such as Bias Education and One-Shot prompting. These results further highlight the importance of bias mitigation strategies in enhancing model reliability. Additional information on the performance of models by bias category and medical specialty across different bias mitigation strategies can be found in Figures H1 and I1 of the appendix, respectively.
@@ 6 Discussion
Our findings are consistent with prior research on medical LLM benchmarking, where proprietary high-resource models outperform open-access models in structured tasks but perform comparably in generative ones. Chen et al. (2025) demonstrated that proprietary high-resource models like GPT-4 and Med-PaLM 2 excel in multiple-choice and factual retrieval tasks, largely due to stronger pretraining on structured datasets and domain-specific knowledge integration. Similarly, Alonso et al. (2024) found that proprietary high-resource models generally achieve higher accuracy in medical question-answering tasks across multiple languages, reinforcing the idea that closed models have better knowledge grounding and factual consistency. Our results support these trends, with Gemini 1.5 Pro and Claude 3.5 Sonnet leading in multiple-choice and fill-in-the-blank tasks, suggesting that API-based models are better suited for clinical decision support and structured question-answering tasks.
Jais underperforms even compared to a random baseline in structured tasks due to severe hallucinations, generating unreliable responses, while Falcon struggles to generalize due to limited exposure to high-quality medical datasets(Penedo et al., 2023b). Figure 3 further illustrates these issues, particularly in the fill-in-the-blank task, where Jais frequently selects invalid options or provides factually inconsistent answers. These results highlight the limitations of open-access models in structured medical applications, where their general-purpose pretraining leads to factual inconsistencies, necessitating domain-specific fine-tuning. Additionally, the relatively lower performance of the Arabic-trained Jais relative to other open-access multilingual models such as Qwen suggests that task specificity may be more critical than language specificity, for Arabic medical tasks, and warrants further exploration in the future.
@@
Figure 4 further illustrates the variability in model performance when exposed to biased content and the effectiveness of different bias mitigation strategies for the top 3 scoring models (GPT-4, Gemini 1.5 Pro, and Claude 3.5 Sonnet-20240620). All models experience an accuracy drop when bias is introduced, with GPT-4 Sonnet showing the most pronounced decline across multiple bias categories. Notably, bias mitigation techniques such as One-Shot and Few-Shot prompting have shown significant improvement in model performance, especially for Gemini 1.5 Pro, which shows the greatest resilience against bias. Yet, no single strategy consistently mitigates bias across all models and categories, underscoring the complexity of bias in medical NLP and the need for further research.
@@
For generative Q&A tasks, however, traditional automatic evaluation metrics such as BERTScore do not fully capture actual model performance. Although GPT-4 and Claude 3.5 produce responses that are highly relevant and contextually accurate as shown in Figure 3, their longer responses result in lower BERTScores compared to ground truth references. This issue has been previously noted by Liu et al. (2025) who found that metrics like ROUGE and BERTScore struggle to effectively assess medical LLMs due to the inherent complexity of medical diagnoses, where multiple treatment options may exist for a single patient. Our findings reinforce this limitation, as models with lower BERTScores sometimes generate high-quality, informative answers that are penalized for verbosity rather than inaccuracy.
To address this limitation, we conducted an additional analysis using GPT-4 as an LLM-based judge, following prior work (Thakur et al., 2025). Each model’s response was rated on a scale of 1 to 5 across four dimensions: similarity to ground truth, relevance/helpfulness, factuality, and safety. This evaluation revealed discrepancies with BERTScore rankings. For example, while Jais had the highest BERTScore (85.7), it received an average GPT score of only 3.2, with models like Deepseek (4.2), Gemini (4.1), and GPT-4 (3.9) outperforming it—particularly in factuality and relevance. Falcon similarly scored well under BERTScore but received the lowest human-aligned rating (1.1), indicating the metric’s failure to capture hallucinations. We also examined potential self-enhancement bias and found none, as GPT-4 rated other models more favorably than itself. These findings, detailed in Appendix J, highlight the limitations of surface-level overlap metrics and underscore the value of more nuanced evaluation techniques such as expert assessments or task-oriented dialogue evaluation for real-world deployment in medical settings.
Ethical Considerations
@@
Given the sensitive nature of healthcare applications, ethical considerations are paramount when developing and deploying LLMs. Our benchmark, MedArabiQ, was carefully constructed from publicly available and anonymized educational resources, ensuring compliance with data privacy standards by excluding personally identifiable patient information. Furthermore, we recognize and discuss several ethical implications associated with clinical deployment of LLMs, including potential risks such as misinformation, unintended amplification of biases (cultural, confirmation, or status quo biases), and limited model interpretability. Addressing these concerns, we emphasize the importance of comprehensive validation through hybrid frameworks combining automated evaluation metrics with clinician expert reviews. We advocate for continuous monitoring, clinician oversight, and clearly defined operational guidelines to mitigate potential harm and biases. By explicitly detailing these ethical considerations, we aim to promote responsible and transparent adoption of LLMs, ultimately contributing to safer and more equitable healthcare solutions.
Limitations
@@
While our study provides a comprehensive evaluation of Arabic medical LLMs, there are areas that warrant further exploration. First, there is a possibility for contamination in the datasets. Although the past medical exams are not available in structured digital format and required an extensive effort for digitization and cleaning, we cannot completely rule out the possibility of contamination. That being said, model performance can be improved highlighting the importance of the benchmark datasets. As for the patient-doctor Q&A data, we explicitly incorporated modifications to test for potential memorization considering that it is sourced from a publicly available dataset. To verify the validity of the dataset, preliminary results were obtained in collaboration with medical students to assess the Q&A dataset for relevance, factuality, complexity, and clarity on a scale from 1 to 5. The average scores (Relevance: 4.99, Factuality: 4.97, Accuracy: 4.88, Clarity: 4.89) provide initial evidence of the dataset’s reliability and utility. Extensive evaluation is required in future work as the scope and size of the benchmark is expanded. A systematic human evaluation, such as expert annotation, would further strengthen the dataset’s credibility, and we plan to incorporate this in future releases.
@@
Furthermore, our benchmarks are in MSA, following standardization practices to ensure consistency. However, this does not account for dialectal variations, which can be problematic in real-world patient-doctor interactions. Future work could explore incorporating dialectal data to enhance model adaptability across diverse Arabic-speaking healthcare contexts. Additionally, while this study focuses on text-based benchmarks as a unimodal foundation, expanding Arabic medical NLP benchmarks to support multimodal inputs—such as medical images and lab results—would be a valuable direction for future research.
@@
Overall, in this work, we introduced the first structured benchmark for evaluating LLMs in Arabic healthcare, addressing a critical gap in Arabic medical NLP. Our benchmark consists of 700 diverse clinical samples, covering both structured medical knowledge assessments and real-world patient-doctor interactions. Beyond Arabic healthcare, our benchmark lays the foundation for developing benchmarks in other medically underserved languages, contributing to the global refinement of medical AI applications.
@@
7 Acknowledgments
This work was supported by the NYUAD Center for Artificial Intelligence and Robotics, funded by Tamkeen under the NYUAD Research Institute Award CG010, the Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104, and the Meem Foundation. The research was carried out on the High Performance Computing resources at New York University Abu Dhabi (Jubail).